Show that the condition for vapor-liquid equilibrium at constant N, T, and P is Gv=GL
Answer #1

The chemical potential of the liquid phase and the vapor phase should be equal at equilibrium. This is the general criterion for phase equlibrium.

\mu^V=\mu^L

However, for a pure component the chemical potential is reduces to the pure component molar Gibbs

free energy,i.e., \mu^V=G^Vand \mu^L=G^L.

So G^V=G^L